# Analyzing PDFs like Binaries: Adversarially Robust PDF Malware Analysis via Intermediate Representation and Language Model

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3744829isbn: 979-8-4007-1525-9/2025/10
@@ 1. Introduction
The Portable Document Format (PDF) is among the most widely used file formats on the web (Johnson, 2021), making it an attractive target for cybercriminals due to its ubiquity and versatility (Palo Alto Networks, 2023; Shah and Saxena, 2024; Bautista, 2023; David Bloxberg, 2024). PDFs now dominate as the most commonly used malicious attachments in phishing campaigns, with nearly 70% of these emails evading network-based defenses and 15% bypassing endpoint security measures (CISA, 2022). Moreover, the increasing prevalence of cloud-based collaboration and remote work has led to the widespread integration of PDF readers within modern browsers, further heightening the threat posed by malicious PDFs. Attackers can exploit vulnerabilities in web applications, such as cross-site scripting (XSS), or browser-specific security flaws to execute malicious code through carefully crafted PDFs (Zhang, 2019; SOCRadar, 2024).
@@
Machine learning (ML) is now extensively applied in various security contexts, including traffic detection, intrusion detection, vulnerability search, and other critical areas. In the face of PDF malware, numerous countermeasures have explored ML-based approaches (Smutz and Stavrou, 2012, 2016; Maiorca et al., 2015; Šrndic and Laskov, 2013; Šrndić and Laskov, 2016; Tong et al., 2019; Chen et al., 2020). However, despite significant progress in PDF malware analysis, several critical challenges remain unresolved.
@@
Firstly, one major challenge is the limited scope of existing feature analysis, which tends to be confined to surface-level inspection. Unlike the advanced feature analysis methods employed in binary code analysis—where researchers strive to extract semantic features from disassembled code and structural features from control flow graphs (CFGs)—PDF malware analysis often remains comparatively rudimentary. Current approaches typically involve computing specific keywords (Smutz and Stavrou, 2012) or analyzing structural paths (Šrndic and Laskov, 2013), which, while useful, fall short of the sophistication needed for gaining deeper insights. While the structure of PDF files differs from that of executable file formats, it presents its own set of unique complexities. Therefore, there is a pressing need for more advanced analysis techniques to effectively address the nuances of PDF malware.
@@
Secondly, the superficial nature of current feature analysis leaves existing PDF malware classifiers highly susceptible to adversarial attacks (Šrndić and Laskov, 2014; Xu et al., 2016; Dang et al., 2017; Maiorca et al., 2019). While some studies (Tong et al., 2019; Chen et al., 2020) have attempted to enhance the adversarial robustness of ML-based classifiers through specialized techniques like adversarial training, they have not substantially advanced feature engineering. Instead, these efforts continue to rely on the simplistic features extracted in earlier work  (Smutz and Stavrou, 2012; Šrndic and Laskov, 2013). Research (Maiorca et al., 2013; Šrndić and Laskov, 2014; Xu et al., 2016) has consistently demonstrated that these features are vulnerable to adversarial manipulation. Additionally, retraining PDF malware classifiers using adversarial samples has significantly compromised their usability, leading to false positive rates (FPR) as high as 85% (Evans et al., 2017).
@@
Lastly, the feature extraction process of existing machine learning-based PDF malware classifiers (Laskov and Šrndić, 2011; Liu et al., 2014; Šrndić and Laskov, 2016; Chen et al., 2020) is heavily dependent on the parsing capabilities of PDF parsers (Carmony et al., 2016; Anantharaman et al., 2023). PDF malware that exploits vulnerabilities in PDF readers often results in so-called ”bad PDFs.” Attackers craft the original bytes of these PDFs to exploit specific vulnerabilities, which frequently results in malformed file formats (Anantharaman et al., 2023). Such malformed PDFs can challenge existing parsers, rendering them unable to correctly process these files, even though the malicious payload can still be successfully executed. Carmony et al. (Carmony et al., 2016) tested multiple parsers on PDF malware datasets and found that each parser failed to correctly process hundreds of samples. Consequently, PDF malware classifiers relying on these parsers cannot extract features from these samples, making it impossible to determine their maliciousness. This limitation is unacceptable in practical applications, where comprehensive detection and analysis of malware are crucial.
@@
To address the second issue, we developed the PDFObj IR representation learning method, PDFObj2Vec, a novel PDF feature engineering approach. We designed three representation learning schemes for this PDFObj2Vec, based on Word2Vec (Mikolov et al., 2013), PV-DM (Le and Mikolov, 2014), and BERT (Devlin et al., 2019). Additionally, we supported PDFObj2Vec with general text embedding models such as CodeT5 (Wang et al., 2021) and text-embedding-3 (OpenAI, 2024), to directly obtain embeddings of PDFObj IR at the ORG node level. We then designed a Graph Isomorphism Network (GIN) to extract structural features at the graph level of the ORG for PDF malware classification.
Graph structures typically exhibit stronger adversarial robustness, making it challenging for attackers to disguise their behavior within such structures (Akoglu et al., 2015). Therefore, our feature engineering approach, which combines semantic and structural features, demonstrates robust performance. This robustness is evidenced by our experimental results, which show strong resilience against various adversarial attacks. While language models have been extensively researched in the context of binary code analysis (Massarelli et al., 2019; Zuo et al., 2019; Chua et al., 2017; Ding et al., 2019; Li et al., 2021), their application to PDF malware analysis has been limited. Our research bridges this gap by applying popular language models, including large language models, to PDF malware analysis. Previous features did not integrate well with these language models, but PDFObj IR demonstrates excellent compatibility, improving the performance of PDF malware analysis tasks.
@@
We applied ORG and PDFObj2Vec to the task of PDF malware classification and implement a GIN-based classifier that achieves well consistent performance on both the baseline and extended datasets. Our most robust classifier attains an accuracy of 99.93% on the baseline dataset and 96.62% on the extended dataset. We also conducted extensive comparative and ablation experiments. The results indicate the effectiveness of PDFObj IR in PDF malware analysis, as our classifiers achieved a 2.2% to 8.9% accuracy improvement on the extended dataset compared to classifiers without PDFObj IR. Furthermore, even when faced with the most powerful realizable adversarial sample attacks, our classifier maintained 100% adversarial robustness with a remarkably low FPR of only 0.07%. This performance is significantly more efficient compared to state-of-the-art PDF malware classifiers (Chen et al., 2020; Tong et al., 2019) with comparable adversarial robustness, whose FPR is 71 times higher than ours. In a nutshell, we make the following key contributions:
@@ 2.1. PDF Basics
PDF Structure PDF is one of the most commonly used document formats on the web (Johnson, 2021), with widespread applications in both personal and business contexts. Figure 1(a) illustrates a typical structure of a PDF file, which consists of the following four parts (ISO, 2008):
@@
  • Body The body of a PDF file comprises various types of objects, forming a collection of objects. The core component of a PDF file is the collection of these objects, also known as COS (Carousel Object System) (Lowagie, 2010) objects.
@@
The Body is the key part of a PDF, containing the primary data in the document. It comprises a series of objects, with each object enclosed by the $\ll$ and $\gg$, demarcated by the obj and endobj. Conceptually, it can be viewed as an object graph, where each object performs specific operations (e.g., displaying text, rendering images, executing code, etc.) (Maiorca et al., 2019). Each object is composed of a series of key-value pairs, which can be represented in the form of a dictionary. For instance, in Figure 1(a), the first object begins with 1 0 obj and ends with endobj. The content of this object is enclosed within $\ll$ $\gg$ and contains four key-value pairs. The keys in the object are the name type, and the values can be any type. For example, The first key in 1 0 obj is /Type, which is a name type with the value of /Catalog, also a name type. The second key, /Outlines, has a value that is an indirect reference type, where R signifies an indirect reference. In a PDF object, values encompass five categories of basic types, as shown in Table 7 of Appendix A. Moreover, values can be composite types, such as arrays and dictionaries, with the basic elements of arrays and dictionaries being the aforementioned basic objects, compression parameters, and other information.
@@ 2.2. Motivation and Insight
In binary analysis, the disassembled code is often scrutinized. For instance, when analyzing a sample in IDA Pro (Hex-Rays, 2025), the software typically presents a CFG of the entry point function after loading the sample. Each basic block corresponds to a disassembled code block, representing the smallest unit of code, as illustrated in Figure 1(c). This raises the question: can we transform the content of a PDF object into a form similar to disassembled code or intermediate code? To explore this, we examined the structure of PDF objects. These objects consist of key-value pairs with fixed data formats and types. By defining a fixed format to describe each object in a manner akin to natural or programming languages, we can convert the object into an intermediate representation. We can then designate the values containing reference relationships as jump instructions. This approach allows us to construct a graph structure that is closely analogous to a CFG, which we refer to as an Object Reference Graph (ORG). Given that we propose converting objects into an intermediate representation, we can leverage powerful language models to process this representation. In the field of program analysis, there has been considerable research using language models to represent disassembled code. We will elaborate on our design philosophy and processing approach in the following sections.
@@ 2.3. PDF Malware Analysis
Currently, research on PDF malware analysis mainly builds upon two prior works: PDFrate (Smutz and Stavrou, 2012) and Hidost (Šrndic and Laskov, 2013). PDFrate utilizes content-based features, extracting specific keyword positions and counts from metadata and content within PDF files. It manually defines 202 features, which are extracted using regular expressions. While these features are more general and not affected by the parsing capabilities of parsers, they remain at a surface level, not delving into the deeper structure of PDFs. Moreover, defining these 202 features requires extensive expert knowledge, and the reliability of these features is not always guaranteed (Šrndić and Laskov, 2014; Tong et al., 2019). Hidost, on the other hand, employs structure-based features by extracting object structural paths from PDFs and using binary counts of these paths as features. It leverages Poppler (Poppler, 2025) to extract hierarchical structural paths and selects $6,087$ paths from a corpus of 9 million as features. Despite the authors’ claims that hierarchical path features are robust (Šrndic and Laskov, 2013; Šrndić and Laskov, 2016), selecting only a portion of the paths from the corpus may result in a lack of path semantics. Xu et al.’s study (Xu et al., 2016) indicates that such features remain vulnerable to adversarial attacks. Furthermore, Hidost relies on Poppler for PDF parsing, which makes feature extraction more susceptible to the parsing capabilities of the parser.
@@
Due to the susceptibility of PDFrate and Hidost to adversarial attacks, researchers have sought to enhance the adversarial robustness of PDF malware classifiers based on their features. Tong et al. (Tong et al., 2019) proposed a PDF malware classifier that leverages conserved feature training, focusing on features derived from PDFrate and Hidost. They identified features closely related to malicious functionalities in PDF execution as conserved features through expert experience. By employing iterative adversarial training, they improved the classifier’s adversarial robustness. However, while Tong et al.’s method enhances robustness against adversarial attacks, it compromises classification performance on regular samples, resulting in a higher FPR of 4.96%.
@@
Additionally, Chen et al. (Chen et al., 2020) proposed a robust training approach based on robust properties targeting the features of Hidost. They defined five categories of robust properties and used symbolic interval analysis to train combinations of different robust properties, resulting in a classifier with adversarial robustness. However, Chen et al.’s method, while achieving adversarial robustness, also sacrifices the classification performance on regular samples, increasing the FPR by 1.78%. Moreover, when facing state-of-the-art unbounded adversarial sample attacks (Maiorca et al., 2013), it only achieved a 50.8% adversarial success rate. Both Tong et al. and Chen et al. attempted to enhance the classifier’s adversarial robustness based on PDFrate and Hidost features, addressing the training issues of the classifier. However, they did not resolve the fundamental problem of insufficiently robust features and the limitations imposed by the parser dependencies in classifiers based on Hidost.
@@ 2.4. Learning-based Embedding
Drawing inspiration from representation learning in binary analysis (Massarelli et al., 2019; Zuo et al., 2019; Chua et al., 2017; Ding et al., 2019), we aim to develop semantic representations for PDFObj IR nodes in the ORG and apply them to downstream tasks such as PDF malware detection. Notably, this type of representation learning remains unexplored in PDF analysis. In binary analysis, approaches like Word2Vec (Mikolov et al., 2013) have been used to learn instruction-level representations by treating each instruction as a word and each function as a document (Massarelli et al., 2019; Zuo et al., 2019). Asm2Vec (Ding et al., 2019) extends this by representing assembly instructions as opcodes and operands, using the PV-DM model (Le and Mikolov, 2014) to learn embeddings. Similarly, PalmTree (Li et al., 2021) treats assembly instructions as sentences, decomposing them into tokens (e.g., opcodes, registers, immediate values) and employing BERT (Devlin et al., 2019) to capture control flow and data dependencies.
In this paper, we applied Word2Vec, PV-DM, and BERT to learn representations for PDFObj IR, generating embeddings for downstream tasks. Additionally, we used general embedding models without pre-training to directly derive embeddings for PDFObj IR. The evaluation of these embedding models in PDF malware classification is detailed in §$\S$6.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; booktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025doi: 10.1145/3719027.3744829isbn: 979-8-4007-1525-9/2025/10
@@ 1. Introduction
The Portable Document Format (PDF) is among the most widely used file formats on the web (pdfa_popularity), making it an attractive target for cybercriminals due to its ubiquity and versatility (2023Unit42Report; macfee_pdf; trustwave_pdf; PDFs_Why). PDFs now dominate as the most commonly used malicious attachments in phishing campaigns, with nearly 70% of these emails evading network-based defenses and 15% bypassing endpoint security measures (Phishing2022cisa). Moreover, the increasing prevalence of cloud-based collaboration and remote work has led to the widespread integration of PDF readers within modern browsers, further heightening the threat posed by malicious PDFs. Attackers can exploit vulnerabilities in web applications, such as cross-site scripting (XSS), or browser-specific security flaws to execute malicious code through carefully crafted PDFs (pdf_nsfocous; SOCRadar2024).
@@
Machine learning (ML) is now extensively applied in various security contexts, including traffic detection, intrusion detection, vulnerability search, and other critical areas. In the face of PDF malware, numerous countermeasures have explored ML-based approaches (smutz2012malicious; smutz2016when; maiorca2015structural; vsrndic2013detection; vsrndic2016hidost; tong2019improving; chen2020training). However, despite significant progress in PDF malware analysis, several critical challenges remain unresolved.
@@
Firstly, one major challenge is the limited scope of existing feature analysis, which tends to be confined to surface-level inspection. Unlike the advanced feature analysis methods employed in binary code analysis—where researchers strive to extract semantic features from disassembled code and structural features from control flow graphs (CFGs)—PDF malware analysis often remains comparatively rudimentary. Current approaches typically involve computing specific keywords (smutz2012malicious) or analyzing structural paths (vsrndic2013detection), which, while useful, fall short of the sophistication needed for gaining deeper insights. While the structure of PDF files differs from that of executable file formats, it presents its own set of unique complexities. Therefore, there is a pressing need for more advanced analysis techniques to effectively address the nuances of PDF malware.
@@
Secondly, the superficial nature of current feature analysis leaves existing PDF malware classifiers highly susceptible to adversarial attacks (vsrndic2014practical; xu2016automatically; dang2017evading; maiorca2019towards). While some studies (tong2019improving; chen2020training) have attempted to enhance the adversarial robustness of ML-based classifiers through specialized techniques like adversarial training, they have not substantially advanced feature engineering. Instead, these efforts continue to rely on the simplistic features extracted in earlier work (smutz2012malicious; vsrndic2013detection). Research (maiorca2013looking; vsrndic2014practical; xu2016automatically) has consistently demonstrated that these features are vulnerable to adversarial manipulation. Additionally, retraining PDF malware classifiers using adversarial samples has significantly compromised their usability, leading to false positive rates (FPR) as high as 85% (AdversarialEvans).
@@
Lastly, the feature extraction process of existing machine learning-based PDF malware classifiers (Laskov2011Static; liu2014detecting; vsrndic2016hidost; chen2020training) is heavily dependent on the parsing capabilities of PDF parsers (carmony2016extract; anantharaman2023polydoc). PDF malware that exploits vulnerabilities in PDF readers often results in so-called ”bad PDFs.” Attackers craft the original bytes of these PDFs to exploit specific vulnerabilities, which frequently results in malformed file formats (anantharaman2023polydoc). Such malformed PDFs can challenge existing parsers, rendering them unable to correctly process these files, even though the malicious payload can still be successfully executed. Carmony et al. (carmony2016extract) tested multiple parsers on PDF malware datasets and found that each parser failed to correctly process hundreds of samples. Consequently, PDF malware classifiers relying on these parsers cannot extract features from these samples, making it impossible to determine their maliciousness. This limitation is unacceptable in practical applications, where comprehensive detection and analysis of malware are crucial.
@@
To address the second issue, we developed the PDFObj IR representation learning method, PDFObj2Vec, a novel PDF feature engineering approach. We designed three representation learning schemes for this PDFObj2Vec, based on Word2Vec (mikolov2013distributed), PV-DM (le2014distributed), and BERT (devlin2019bert). Additionally, we supported PDFObj2Vec with general text embedding models such as CodeT5 (wang-etal-2021-codet5) and text-embedding-3 (openai_embedding), to directly obtain embeddings of PDFObj IR at the ORG node level. We then designed a Graph Isomorphism Network (GIN) to extract structural features at the graph level of the ORG for PDF malware classification.
Graph structures typically exhibit stronger adversarial robustness, making it challenging for attackers to disguise their behavior within such structures (akoglu2015graph). Therefore, our feature engineering approach, which combines semantic and structural features, demonstrates robust performance. This robustness is evidenced by our experimental results, which show strong resilience against various adversarial attacks. While language models have been extensively researched in the context of binary code analysis (massarelli2019safe; zuo2018neural; chua2017neural; ding2019asm2vec; li2021palmtree), their application to PDF malware analysis has been limited. Our research bridges this gap by applying popular language models, including large language models, to PDF malware analysis. Previous features did not integrate well with these language models, but PDFObj IR demonstrates excellent compatibility, improving the performance of PDF malware analysis tasks.
@@
We applied ORG and PDFObj2Vec to the task of PDF malware classification and implement a GIN-based classifier that achieves well consistent performance on both the baseline and extended datasets. Our most robust classifier attains an accuracy of 99.93% on the baseline dataset and 96.62% on the extended dataset. We also conducted extensive comparative and ablation experiments. The results indicate the effectiveness of PDFObj IR in PDF malware analysis, as our classifiers achieved a 2.2% to 8.9% accuracy improvement on the extended dataset compared to classifiers without PDFObj IR. Furthermore, even when faced with the most powerful realizable adversarial sample attacks, our classifier maintained 100% adversarial robustness with a remarkably low FPR of only 0.07%. This performance is significantly more efficient compared to state-of-the-art PDF malware classifiers (chen2020training; tong2019improving) with comparable adversarial robustness, whose FPR is 71 times higher than ours. In a nutshell, we make the following key contributions:
@@ 2.1. PDF Basics
PDF Structure PDF is one of the most commonly used document formats on the web (pdfa_popularity), with widespread applications in both personal and business contexts. Figure 1(a) illustrates a typical structure of a PDF file, which consists of the following four parts (pdf_standard):
@@
  • Body The body of a PDF file comprises various types of objects, forming a collection of objects. The core component of a PDF file is the collection of these objects, also known as COS (Carousel Object System) (lowagie2010itext) objects.
@@
The Body is the key part of a PDF, containing the primary data in the document. It comprises a series of objects, with each object enclosed by the $\ll$ and $\gg$, demarcated by the obj and endobj. Conceptually, it can be viewed as an object graph, where each object performs specific operations (e.g., displaying text, rendering images, executing code, etc.) (maiorca2019towards). Each object is composed of a series of key-value pairs, which can be represented in the form of a dictionary. For instance, in Figure 1(a), the first object begins with 1 0 obj and ends with endobj. The content of this object is enclosed within $\ll$ $\gg$ and contains four key-value pairs. The keys in the object are the name type, and the values can be any type. For example, The first key in 1 0 obj is /Type, which is a name type with the value of /Catalog, also a name type. The second key, /Outlines, has a value that is an indirect reference type, where R signifies an indirect reference. In a PDF object, values encompass five categories of basic types, as shown in Table 7 of Appendix A. Moreover, values can be composite types, such as arrays and dictionaries, with the basic elements of arrays and dictionaries being the aforementioned basic objects, compression parameters, and other information.
@@ 2.2. Motivation and Insight
In binary analysis, the disassembled code is often scrutinized. For instance, when analyzing a sample in IDA Pro (idapro), the software typically presents a CFG of the entry point function after loading the sample. Each basic block corresponds to a disassembled code block, representing the smallest unit of code, as illustrated in Figure 1(c). This raises the question: can we transform the content of a PDF object into a form similar to disassembled code or intermediate code? To explore this, we examined the structure of PDF objects. These objects consist of key-value pairs with fixed data formats and types. By defining a fixed format to describe each object in a manner akin to natural or programming languages, we can convert the object into an intermediate representation. We can then designate the values containing reference relationships as jump instructions. This approach allows us to construct a graph structure that is closely analogous to a CFG, which we refer to as an Object Reference Graph (ORG). Given that we propose converting objects into an intermediate representation, we can leverage powerful language models to process this representation. In the field of program analysis, there has been considerable research using language models to represent disassembled code. We will elaborate on our design philosophy and processing approach in the following sections.
@@ 2.3. PDF Malware Analysis
Currently, research on PDF malware analysis mainly builds upon two prior works: PDFrate (smutz2012malicious) and Hidost (vsrndic2013detection). PDFrate utilizes content-based features, extracting specific keyword positions and counts from metadata and content within PDF files. It manually defines 202 features, which are extracted using regular expressions. While these features are more general and not affected by the parsing capabilities of parsers, they remain at a surface level, not delving into the deeper structure of PDFs. Moreover, defining these 202 features requires extensive expert knowledge, and the reliability of these features is not always guaranteed (vsrndic2014practical; tong2019improving). Hidost, on the other hand, employs structure-based features by extracting object structural paths from PDFs and using binary counts of these paths as features. It leverages Poppler (Poppler55) to extract hierarchical structural paths and selects $6,087$ paths from a corpus of 9 million as features. Despite the authors’ claims that hierarchical path features are robust (vsrndic2013detection; vsrndic2016hidost), selecting only a portion of the paths from the corpus may result in a lack of path semantics. Xu et al.’s study (xu2016automatically) indicates that such features remain vulnerable to adversarial attacks. Furthermore, Hidost relies on Poppler for PDF parsing, which makes feature extraction more susceptible to the parsing capabilities of the parser.
@@
Due to the susceptibility of PDFrate and Hidost to adversarial attacks, researchers have sought to enhance the adversarial robustness of PDF malware classifiers based on their features. Tong et al. (tong2019improving) proposed a PDF malware classifier that leverages conserved feature training, focusing on features derived from PDFrate and Hidost. They identified features closely related to malicious functionalities in PDF execution as conserved features through expert experience. By employing iterative adversarial training, they improved the classifier’s adversarial robustness. However, while Tong et al.’s method enhances robustness against adversarial attacks, it compromises classification performance on regular samples, resulting in a higher FPR of 4.96%.
@@
Additionally, Chen et al. (chen2020training) proposed a robust training approach based on robust properties targeting the features of Hidost. They defined five categories of robust properties and used symbolic interval analysis to train combinations of different robust properties, resulting in a classifier with adversarial robustness. However, Chen et al.’s method, while achieving adversarial robustness, also sacrifices the classification performance on regular samples, increasing the FPR by 1.78%. Moreover, when facing state-of-the-art unbounded adversarial sample attacks (maiorca2013looking), it only achieved a 50.8% adversarial success rate. Both Tong et al. and Chen et al. attempted to enhance the classifier’s adversarial robustness based on PDFrate and Hidost features, addressing the training issues of the classifier. However, they did not resolve the fundamental problem of insufficiently robust features and the limitations imposed by the parser dependencies in classifiers based on Hidost.
@@ 2.4. Learning-based Embedding
Drawing inspiration from representation learning in binary analysis (massarelli2019safe; zuo2018neural; chua2017neural; ding2019asm2vec), we aim to develop semantic representations for PDFObj IR nodes in the ORG and apply them to downstream tasks such as PDF malware detection. Notably, this type of representation learning remains unexplored in PDF analysis. In binary analysis, approaches like Word2Vec (mikolov2013distributed) have been used to learn instruction-level representations by treating each instruction as a word and each function as a document (massarelli2019safe; zuo2018neural). Asm2Vec (ding2019asm2vec) extends this by representing assembly instructions as opcodes and operands, using the PV-DM model (le2014distributed) to learn embeddings. Similarly, PalmTree (li2021palmtree) treats assembly instructions as sentences, decomposing them into tokens (e.g., opcodes, registers, immediate values) and employing BERT (devlin2019bert) to capture control flow and data dependencies.
In this paper, we applied Word2Vec, PV-DM, and BERT to learn representations for PDFObj IR, generating embeddings for downstream tasks. Additionally, we used general embedding models without pre-training to directly derive embeddings for PDFObj IR. The evaluation of these embedding models in PDF malware classification is detailed in $\S$6.